How many litres of pure alcohol must be added to 20 litres of a 50% alcohol concentration in order to raise it to a 60% alcohol concentration?
Answer by richwmiller(17219)

x+20/2=.6(x+20)
x+10=.6(x+20)
.4x=12-10
.4x=2
4/10x=2
4x=20
x=5 litres of pure alcohol to be added
check
5+10=.6(25)
15=15
ok